# 1]DAMO Academy, Alibaba Group 2]Hupan Lab

### v1

@@ Abstract
Recent advancements in Large Vision-Language Models (LVLMs) have significantly enhanced their ability to integrate visual and linguistic information, achieving near-human proficiency in tasks like object recognition, captioning, and visual question answering. However, current benchmarks typically focus on knowledge-centric evaluations that assess domain-specific expertise, often neglecting the core ability to reason about fundamental mathematical elements and visual concepts. We identify a gap in evaluating elementary-level math problems, which rely on explicit visual dependencies-requiring models to discern, integrate, and reason across multiple images while incorporating commonsense knowledge, all of which are crucial for advancing toward broader AGI capabilities. To address this gap, we introduce VCBench, a comprehensive benchmark for multimodal mathematical reasoning with explicit visual dependencies. VCBench includes 1,720 problems across six cognitive domains, featuring 6,697 images (averaging 3.9 per question) to ensure multi-image reasoning. We evaluate 26 state-of-the-art LVLMs on VCBench, revealing substantial performance disparities, with even the top models unable to exceed 50% accuracy. Our findings highlight the ongoing challenges in visual-mathematical integration and suggest avenues for future LVLM advancements.
@@ 1 Introduction
While many visual mathematics benchmarks, such as MathVista Lu et al. (2023) and MathVision Wang et al. (2024a), focus on knowledge-centric evaluations that assess domain-specific mathematical or geometric expertise, they often fail to evaluate a model’s core ability to perceive and reason about fundamental mathematical elements and visual concepts. Moreover, these knowledge-centric evaluations are easily influenced by the pre-existing knowledge embedded in large language models, which may obscure true reasoning capabilities. To advance towards Artificial General Intelligence (AGI), a more holistic approach to multi-modal reasoning is needed-one that goes beyond task-specific benchmarks and better captures generalizable cognitive abilities.
In this context, we identify a gap in the evaluation of models on elementary-level math problems Cobbe et al. (2021a); Wei et al. (2023b). These problems, typically at the elementary school level, do not require complex mathematical or geometric reasoning but rely heavily on explicit visual dependencies-the ability to discern and integrate visual features across images and understand how different visual elements relate to one another to solve problems. This mirrors the cognitive development of children, who, at a young age, rely on similar skills to solve problems despite not yet possessing advanced reasoning abilities. Understanding and modeling this form of reasoning is crucial, as it represents a fundamental cognitive ability essential for advancing toward broader AGI capabilities.
@@
Unlike existing benchmarks that focus on knowledge-centric evaluations, we emphasize vision-centric assessments. VCBench targets problems that do not require specialized knowledge but instead rely on the common perceptual reasoning of mathematical images and concepts. This approach aligns with the way children learn-first mastering visual reasoning and only later acquiring domain-specific knowledge.
@@
Our benchmark provides a holistic evaluation of various visual reasoning capabilities, such as temporal reasoning, spatial understanding, and pattern recognition. While these tasks may seem simple to children, they represent fundamental reasoning abilities that LVLMs often struggle with. Our experiments demonstrate that tasks considered easy for children-such as identifying time sequences or spatial relationships-prove challenging for state-of-the-art LVLMs, highlighting the gaps in current multimodal reasoning capabilities.
@@ 4.3 Results of Math Specialist Models
The Math Specialist models, which include G-LLaVA-13B, G-LLaVA-7B, and MathLlava, exhibit relatively low overall performance, with average scores ranging from 25.47 to 29.30. Notably, G-LLaVA-13B records the lowest score at 25.47, while MathLlava achieves a slightly higher average of 29.30. Although these models are designed with a focus on mathematical reasoning, their performance across diverse tasks-such as time and calendar, spatial reasoning, and geometric challenges-remains inconsistent. For instance, while G-LLaVA-7B reaches a perfect score (100.00) on the Calendar sub-task, its scores in other categories, such as Clock and certain geometry-related tasks, are considerably lower.
Furthermore, the results indicate that these Math Specialist models struggle to match the performance of their general-purpose counterparts. Despite showing some strengths-for example, MathLlava scoring 34.93 on the Cube task-these models fall short on several key aspects, including Clock, Location, and reasoning tasks. This pattern underscores the challenge of integrating specialized mathematical capabilities with the broader spectrum of visual understanding.
@@ 6 Conclusion
This paper introduces VCBench-a comprehensive evaluation framework designed to assess multimodal mathematical reasoning with explicit visual dependency. By addressing the limitations of existing datasets in multi-image integration and cross-modal relational reasoning, our benchmark provides a detailed analysis of 26 state-of-the-art LVLMs across six cognitive domains and 17 task categories. The evaluation reveals significant performance disparities, particularly in areas such as multi-step instruction following, basic visual perception, cross-image consistency, and vulnerability to visual hallucinations.

### v2

@@ Abstract
Recent advancements in Large Vision-Language Models (LVLMs) have significantly enhanced their ability to integrate visual and linguistic information, achieving near-human proficiency in tasks like object recognition, captioning, and visual question answering. However, current benchmarks typically focus on knowledge-centric evaluations that assess domain-specific expertise, often neglecting the core ability to reason about fundamental mathematical elements and visual concepts. We identify a gap in evaluating elementary-level math problems, which rely on explicit visual dependencies-requiring models to discern, integrate, and reason across multiple images while incorporating commonsense knowledge, all o which are crucial for advancing toward broader AGI capabilities. To address this gap, we introduce VCBench, a comprehensive benchmark for multimodal mathematical reasoning with explicit visual dependencies. VCBench includes 1,720 problems across six cognitive domains, featuring 6,697 images (averaging 3.9 per question) to ensure multi-image reasoning. We evaluate 26 state-of-the-art LVLMs on VCBench, revealing substantial performance disparities, with even the top models unable to exceed 50% accuracy. Our findings highlight the ongoing challenges in visual-mathematical integration and suggest avenues for future LVLM advancements.
@@ 1 Introduction
While many visual mathematics benchmarks, such as MathVista Lu et al. (2023) and MathVision Wang et al. (2024a), focus on knowledge-centric evaluations that assess domain-specific mathematical or geometric expertise, they often fail to evaluate a model’s core ability to perceive and reason about fundamental mathematical elements and visual concepts. Moreover, these knowledge-centric evaluations are easily influenced by the pre-existing knowledge embedded in large language models, which may obscure true reasoning capabilities. To advance towards Artificial General Intelligence (AGI), a more holistic approach to multi-modal reasoning is needed—one that goes beyond task-specific benchmarks and better captures generalizable cognitive abilities.
In this context, we identify a gap in the evaluation of models on elementary-level math problems Cobbe et al. (2021a); Wei et al. (2023b). These problems, typically at the elementary school level, do not require complex mathematical or geometric reasoning but rely heavily on explicit visual dependencies—the ability to discern and integrate visual features across images and understand how different visual elements relate to one another to solve problems. This mirrors the cognitive development of children, who, at a young age, rely on similar skills to solve problems despite not yet possessing advanced reasoning abilities. Understanding and modeling this form of reasoning is crucial, as it represents a fundamental cognitive ability essential for advancing toward broader AGI capabilities.
@@
Unlike existing benchmarks that focus on knowledge-centric evaluations, we emphasize vision-centric assessments. VCBench targets problems that do not require specialized knowledge but instead rely on the common perceptual reasoning of mathematical images and concepts. This approach aligns with the way children learn—first mastering visual reasoning and only later acquiring domain-specific knowledge.
@@
Our benchmark provides a holistic evaluation of various visual reasoning capabilities, such as temporal reasoning, spatial understanding, and pattern recognition. While these tasks may seem simple to children, they represent fundamental reasoning abilities that LVLMs often struggle with. Our experiments demonstrate that tasks considered easy for children—such as identifying time sequences or spatial relationships—prove challenging for state-of-the-art LVLMs, highlighting the gaps in current multimodal reasoning capabilities.
@@ 4.3 Results of Math Specialist Models
The Math Specialist models, which include G-LLaVA-13B, G-LLaVA-7B, and MathLlava, exhibit relatively low overall performance, with average scores ranging from 25.47 to 29.30. Notably, G-LLaVA-13B records the lowest score at 25.47, while MathLlava achieves a slightly higher average of 29.30. Although these models are designed with a focus on mathematical reasoning, their performance across diverse tasks—such as time and calendar, spatial reasoning, and geometric challenges—remains inconsistent. For instance, while G-LLaVA-7B reaches a perfect score (100.00) on the Calendar sub-task, its scores in other categories, such as Clock and certain geometry-related tasks, are considerably lower.
Furthermore, the results indicate that these Math Specialist models struggle to match the performance of their general-purpose counterparts. Despite showing some strengths—for example, MathLlava scoring 34.93 on the Cube task—these models fall short on several key aspects, including Clock, Location, and reasoning tasks. This pattern underscores the challenge of integrating specialized mathematical capabilities with the broader spectrum of visual understanding.
@@ 6 Conclusion
This paper introduces VCBench—a comprehensive evaluation framework designed to assess multimodal mathematical reasoning with explicit visual dependency. By addressing the limitations of existing datasets in multi-image integration and cross-modal relational reasoning, our benchmark provides a detailed analysis of 26 state-of-the-art LVLMs across six cognitive domains and 17 task categories. The evaluation reveals significant performance disparities, particularly in areas such as multi-step instruction following, basic visual perception, cross-image consistency, and vulnerability to visual hallucinations.
Acknowledgement
The LaTeX template is built upon Meta’s original template.